OpenStudy (anonymous):

Solve the equation. |x + 7| = 12 {–19, –5} {–19, 5} {–5, 19} {5, 19}

OpenStudy (zzr0ck3r):

\(\lvert a\rvert +b \implies a=b \text { or } a=-b\)

OpenStudy (anonymous):

I dont get that?

OpenStudy (zzr0ck3r):

\(|x+7|=12\) implies \(x+7=12\) or \(x+7=-12\)